# Generalized Bayesian Inference for Causal Effects using Covariate Balancing Procedure

### v2

@@ 4.1 Data-generating Mechanism
Next, we introduce the assignment mechanism of the treatment value $A_{i}$; that is, the true propensity score $e(\bm{X}_{i})={\rm Pr}\left(A=1|\bm{X}_{i}\right)=expit\left\{h_{A}\left(\bm{X}_{i}\right)\right\}$. Here, we considered six scenarios: “Small complexity and good overlap,” “Small complexity and poor overlap,” “Moderate complexity and good overlap,” “Moderate complexity and poor overlap,” “Large complexity and good overlap,” and “Large complexity and poor overlap.” For each situation, the linear predictor $h_{A}\left(\bm{X}_{i}\right)$ of the true propensity score model was set as follows:
@@
For more details, see Orihara et al.[24].
Finally, we introduce the model for the outcome $Y_{i}$. This simulation considered the binary outcome and logistic regression model ${\rm Pr}\left(Y=1|\bm{X}_{i}\right)=expit\left\{h_{Y}\left(\bm{X}_{i}\right)\right\}$ where
[…]
Under these settings, the true causal effects (i.e., ATE) is $0.152$ for all propensity score situations (a)(f).
4.2 Estimating Methods
We considered four methods for estimating the propensity score: ordinary logistic regression, covariate balancing propensity score (CBPS[16]), regularized calibration method (RCAL[31]), and the proposed Bayesian method.
To estimate the propensity score, we consider two situations: “Confounders” and “Predictors.” For the former situation, we use only the confounders to estimate the propensity score; $\left(X_{i1},X_{i2},X_{i3},X_{i4}\right)$. Thus, we can at least adjust for confounder bias. In the latter situation, we use the predictors of the true propensity score; $\left(X_{i1},X_{i2},\dots,X_{i7}\right)$. Thus, there is a chance to estimate an unbiased propensity score compared to the “Confounders” situations. However, note that adjusting for instrumental variables that are only related to the exposure variable ($X_{i5}$, $X_{i6}$, and $X_{i7}$) can result in unstable causal effect estimates[4].
As mentioned in Section 2, the CBPS is a robust parametric propensity score estimation method that considers model misspecification. In this simulation experiment, we only considered the first-order moment aligning with previous research[28]. Note that other methods also use the same moment condition.
For RCAL method proposed in Tan[31], it is necessary to decide on the tuning parameter $\lambda$, as described in Section 2 (see (2.1)). Specifically, to determine this parameter, we consider two approaches: making a deterministic decision or using Cross-Validation (CV). In the former approach, we prepare some values of $\lambda$ in advance. In the latter approach, CV is conducted with the loss function related to the propensity score (2.2). To implement RCAL, we employ the glm.regu and ate.ipw functions from the RCAL package in R.
The proposed method, called Bayesian RCAL (BRCAL), is based on the estimating procedure that includes estimating the learning rate $\omega$ by PCIC, as explained in Section 3. The program files for implementing our proposed method is stored on the author’s GitHub page.
4.3 Performance Metrics
We evaluated the various methods based on bias, root mean squared error (RMSE), coverage probability (CP), average length of Confidence (Credible) Interval (AvL), bias and RMSE ratio (BR/RR) compared to the proposed method, and boxplot of estimated ATE from 10,000 iterations. The bias and RMSE were calculated as follows:
$bias=\bar{\hat{\tau}}-\tau_{0},\ \ \ RMSE=\sqrt{\frac{1}{10000}\sum_{k=1}^{10000}\left(\hat{\tau}_{k}-\tau_{0}\right)^{2}},$
where $\bar{\hat{\tau}}=\frac{1}{10000}\sum_{k=1}^{10000}\hat{\theta}_{k}$, $\hat{\tau}_{k}$ is the estimate of each estimator and iteration, and $\tau_{0}$ ($=0.152$) is the true value of the ATE. The CP refers to the proportion of cases where the confidence interval or credible interval includes $\tau_{0}$. The AvL was calculated as the average length of the Confidence (Credible) Interval over 10,000–time iterations. Bias and RMSE ratio (BR/RR) are calculated as the ratio compared to the bias/RMSE of our proposed method.
4.4 Simulation Results
The results are summarized in Figures 1 and A.1 and Table A.1A.6 in Appendix A. Note that for Scenario (f), results are summarized only in tables to ensure visibility of figures. In the main manuscript, we only summarize the results under “Confounders” situations.
The ordinary logistic regression exhibits small bias in scenarios (a) and (b) since the true propensity score model is correctly specified. However, in scenarios (c)(e), where the true propensity score model cannot be specified, BRCAL yields attractive results for the bias compared to other methods, especially in small samples. CBPS performs similarly to BRCAL, particularly in large samples. RCAL also performs well in some scenarios, but exhibits unstable bias results in small samples.
All methods exhibit nearly identical RMSE values across all scenarios. If anything, RCAL performs slightly better in small samples.
The 95% coverage probability (CP) of BRCAL achieves the nominal level and is higher compared to other methods, for almost all scenarios. In scenario (e), BRCAL slightly underestimates the 95% CI. In small samples, all methods except for BRCAL underestimate the 95% CI in all scenarios. Notably, RCAL estimates comparatively narrow CIs in scenarios (b) and (e). However, this tendency is observed in other methods as well.
BRCAL, our proposed method, is considered the Bayesian version of RCAL. However, BRCAL’s performance is more akin to that of CBPS, especially in large samples. In small samples, BRCAL performs well in terms of bias and 95% CP compared to CBPS and RCAL.
To summarize, our proposed method demonstrates greater, or at least equal performance compared to other methods. Notably, it yields more accurate causal effect estimates and reliable inference results in small samples.
Figure 1: Summary of estimated causal effects: Bias, root mean squared error (RMSE), and 95% coverage probability (95%CP) in 10,000 iterations were summarized under “Covariate” situation. The left panel: $n=200$; the center panel: $n=500$; The right panel: $n=1000$. RCAL (CV) was used for the CV described in Section 4.2. BRCAL (PCIC) was used for the PCIC to select the learning rate $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
5 Illustrative Real Data Examples
In this section, we provide an illustrative example of a real data analysis conducted on the cohort study for male civil servants in London (Whitehall dataset)[26]. Specifically, we examined the causal effects of smoking on all-cause mortality. To estimate these effects, we considered three confounders: age, systolic blood pressure, and job grade. To account for the binary exposure, we coded the original data with $A_{i}=0$ for non-smokers, and $A_{i}=1$ otherwise. The study included 17,260 subjects, of which 7,157 (41.5%) were smokers. There were also 1,670 deaths (9.7%).
The summary results of the causal effects for each method are as follows: The crude mean difference is $6.86(5.97,7.75)$ (point estimate (95% CI)), the CBPS result is $5.62(5.58,5.68)$, and the RCAL result is $5.71(4.81,6.62)$. The result of our proposed method is $5.65(4.19,7.12)$ (posterior mean (95% CI)). The point estimates are similar for CBPS, RCAL, and BRCAL. The 95% CI for BRCAL is the widest among the three methods. Given that there are 17,260 subjects, the CI is considered too wide. However, the CI is expected to provide conservative results, based on the findings from the simulation experiments.
6 Discussions and Conclusions
In this manuscript, we introduce a novel Bayesian-based IPW estimator that incorporates a covariate balancing condition. Utilizing the General Bayes framework, which is gaining interest in Bayesian analysis, our proposed method does not require consideration of the full likelihood. Accordingly, it only necessitates minimal moment conditions, as is common in causal inference contexts. Our approach demonstrates reasonable performance, especially in small sample scenarios, as shown by various simulations and real data analyses.
To the best of our knowledge, this is the first method that applies the General Bayes framework to causal inference with a covariate balancing condition. As described in Section 3, the learning rate $\omega$ is estimated by PCIC. This method proves effective, as evidenced by the 95% CI aligning well with the nominal probability. However, relying on PCIC is an ad-hoc strategy lacking theoretical justification. In future work, we aim to develop a consensus method that is theoretically grounded.
In our proposed framework, the posterior sampling scheme is conducted in two stages. First, parameters related to the posterior of the propensity score model are sampled, followed by parameters for the IPW estimator. This approach corresponds to the “two-step inference” known in the Bayesian context [20, 29], which allows us to estimate causal effects as in standard causal inference contexts [17]. The inference is justified from a Bayesian perspective [20], provided the second step is likelihood-based. Bissiri et al. [3] note that the general Bayesian procedure ensures valid Bayesian inference by meeting Coherence and Rationality [2]. Future work should explore further justification when applying the general Bayesian framework for the second step.
For future work, applying our approach to sequential treatment settings is crucial for estimating the ATE in situations with variable doses. While some methods incorporating covariate balancing conditions have been proposed [15, 1], a loss function akin to that described by Tan [31] has yet to be developed. Going forward, deriving such a loss function and considering a Bayesian estimation strategy following the same approach as presented in this manuscript will be essential. In sequential treatment settings, concerns about missing data, particularly in scenarios with extended follow-up periods, arise. In Appendix B, we consider the situations where one covariate is missing in point treatment scenarios. However, these methods cannot be directly applied to sequential settings. Xu et al. [35] introduce the Bayesian Additive Regression Tree method for imputing missing data as a Bayesian approach. This method, which is potentially adaptable to our proposed and future methodologies, could enable the construction of a comprehensive Bayesian method that effectively handles missing covariate data.
Acknowledgements: All the authors accept responsibility for the content of this manuscript and approve its submission.
Data availability statement: The data used in the real data analysis cannot be made available for general use. The simulation analysis programs are available at the following URL:
  • https://github.com/t-momozaki/BRCAL
Appendix A Simulation Results
@@
BRCAL and CBPS exhibit almost the same value of bias in almost all scenarios. In contrast, RCAL exhibits large, unstable bias compared to other methods. The results of all methods become consistent in large samples.
Table A.1: Summary of estimated causal effects under the scenario (a): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each situation (in “Covariate” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning rate $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Covariate Method
@@
Confounders Logit 0.302 5.587 0.942 0.214 0.888 1.006 0.024 3.479 0.951 0.136 0.490 1.002 0.057 2.487 0.946 0.096 0.826 1.001
CBPS 0.283 5.585 0.941 0.214 0.832 1.005 0.016 3.480 0.951 0.136 0.327 1.002 0.050 2.487 0.947 0.096 0.725 1.001
RCAL CV 0.696 5.444 0.943 0.208 2.047 0.980 0.170 3.455 0.945 0.132 3.469 0.995 0.122 2.482 0.939 0.093 1.768 0.999
$\lambda=0.01$ 0.601 5.503 0.941 0.208 0.323 3.419 0.947 0.132 0.361 2.444 0.944 0.093
$\lambda=0.05$ 0.967 5.254 0.949 0.207 0.994 3.276 0.954 0.131 1.160 2.345 0.951 0.092
$\lambda=0.10$ 0.762 5.183 0.950 0.206 0.502 3.249 0.954 0.130 0.445 2.327 0.952 0.092
BRCAL PCIC 0.340 5.555 0.961 0.256 Ref. Ref. 0.049 3.473 0.964 0.146 Ref. Ref. 0.069 2.484 0.962 0.104 Ref. Ref.
$\omega=0.2$ 0.427 5.423 1.000 0.619 0.205 3.432 1.000 0.392 0.184 2.471 1.000 0.277
$\omega=0.5$ 0.426 5.511 0.999 0.392 0.099 3.459 1.000 0.248 0.103 2.481 1.000 0.175
$\omega=1.0$ 0.350 5.548 0.988 0.278 0.064 3.470 0.988 0.175 0.079 2.484 0.987 0.124
$\omega=1.5$ 0.337 5.559 0.957 0.227 0.048 3.473 0.961 0.143 0.066 2.485 0.958 0.101
Predictors Logit 0.362 6.454 0.931 0.232 1.003 1.022 0.395 3.977 0.941 0.150 1.079 1.010 0.189 2.752 0.951 0.107 1.212 1.007
CBPS 0.290 6.372 0.927 0.230 0.803 1.010 0.311 3.953 0.938 0.148 0.850 1.004 0.191 2.740 0.949 0.106 1.224 1.002
RCAL CV 0.953 5.799 0.926 0.208 2.640 0.919 0.860 3.846 0.913 0.132 2.350 0.977 0.117 2.708 0.915 0.093 0.750 0.990
$\lambda=0.01$ 1.288 6.654 0.896 0.208 0.834 3.872 0.912 0.132 0.196 2.651 0.922 0.093
$\lambda=0.05$ 0.972 5.595 0.936 0.208 1.373 3.461 0.941 0.131 1.036 2.408 0.946 0.093
$\lambda=0.10$ 0.771 5.342 0.943 0.207 0.877 3.348 0.946 0.130 0.295 2.347 0.949 0.092
BRCAL PCIC 0.361 6.312 0.932 0.231 Ref. Ref. 0.366 3.937 0.934 0.145 Ref. Ref. 0.156 2.734 0.939 0.103 Ref. Ref.
$\omega=0.2$ 0.652 5.993 1.000 0.621 0.628 3.848 1.000 0.392 0.012 2.701 1.000 0.277
$\omega=0.5$ 0.459 6.205 0.998 0.393 0.452 3.910 0.998 0.248 0.108 2.725 0.999 0.175
$\omega=1.0$ 0.389 6.285 0.974 0.278 0.386 3.931 0.974 0.175 0.144 2.732 0.975 0.124
$\omega=1.5$ 0.357 6.314 0.929 0.227 0.362 3.938 0.932 0.143 0.157 2.734 0.937 0.101
@@
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[16]; RCAL: Calibrating method proposed by Tan[31]; BRCAL: Proposed method.
Table A.2: Summary of estimated causal effects under the scenario (b): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each situation (in “Covariate” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Covariate Method

### v3

@@ 4.1 Data-generating Mechanism
Next, we introduce the assignment mechanism of the treatment value $A_{i}$; that is, the true propensity score $e(\bm{X}_{i})={\rm Pr}\left(A=1\mid\bm{X}_{i}\right)=expit\left\{h_{A}\left(\bm{X}_{i}\right)\right\}$. Here, we considered six scenarios by changing the the linear predictor $h_{A}\left(\bm{X}_{i}\right)$; specifically, (a): “Small complexity and good overlap,” (b): “Small complexity and poor overlap,” (c): “Moderate complexity and good overlap,” (d): “Moderate complexity and poor overlap,” (e): “Large complexity and good overlap,” and (f): “Large complexity and poor overlap.” For more details, see Appendix B and Orihara et al.[27].
Finally, we introduce the model for the outcome $Y_{i}$. This simulation considered the binary outcome and logistic regression model ${\rm Pr}\left(Y=1\mid\bm{X}_{i}\right)=expit\left\{h_{Y}\left(\bm{X}_{i}\right)\right\}$ where
[…]
Under these settings, the true causal effects (i.e., ATE) is approximately $0.152$ for all propensity score situations (a)(f).
4.2 Estimating Methods
We considered four methods for estimating the propensity score: ordinary logistic regression, covariate balancing propensity score (CBPS[4]), regularized calibration method (RCAL[10]), a Bayesian method using a Logistic regression-based Propensity Score (BLPS) proposed by Saarela et al. [Sa2015], and the proposed Bayesian method explained in Section 3.
To estimate the propensity score, we consider two situations: “Confounders” and “Predictors.” For the former situation, we use only the confounders to estimate the propensity score; $g(\bm{X}_{i})=\left(X_{i1},X_{i2},X_{i3},X_{i4}\right)$. Thus, we can at least adjust for confounder bias. In the latter situation, we use the predictors of the true propensity score; $g(\bm{X}_{i})=\left(X_{i1},X_{i2},\dots,X_{i7}\right)$. Thus, there is a chance to estimate an unbiased propensity score compared to the “Confounders” situations. However, note that adjusting for variables that are only related to the exposure variable ($X_{i5}$, $X_{i6}$, and $X_{i7}$) can yield an increase in the variance of causal effect estimators[28].
As mentioned in Section 2, the CBPS is a robust parametric propensity score estimation method that considers model misspecification. In this simulation experiment, we only considered the first-order moment aligning with previous research[26]. Note that other methods also use the same moment condition, and the covariates are standardized.
For the RCAL method proposed in Tan[10], it is necessary to decide on the tuning parameter $\lambda$, as described in Section 2 (see (2.1)). Specifically, to determine this parameter, we consider two approaches: making a deterministic decision or using Cross-Validation (CV). In the former approach, we prepare some values of $\lambda$ in advance. In the latter approach, CV is conducted with the loss function related to the propensity score (2.2). To implement the RCAL, we employ the glm.regu and ate.ipw functions from the RCAL package in R.
For the BLPS method proposed by Saarela et al. [Sa2015], the Bayesian bootstrap is used to estimate the ATE. While their method can be applied to sequential treatment settings, we used it only for a point treatment setting.
The proposed method, called Bayesian RCAL (BRCAL), is based on an estimation procedure that includes determining the learning rate $\omega$ using PCIC, as explained in Section 3. Prior distributions of parameters are $\lambda\sim Gam(0.01,0.1)$, $\bm{\alpha}\sim DoubleExponential(0,1/\lambda)$, and $\theta_{k}\sim N(0,100^{2})$ ($k=0,1$). The specific sampling schemes are described in Section 3.2. The program files for implementing BLPS and our proposed method are stored on the author’s GitHub page.
4.3 Performance Metrics
We evaluated the various methods based on bias, root mean squared error (RMSE), coverage probability (CP), average length of Confidence (Credible) Interval (AvL), bias and RMSE ratio (BR/RR) compared to the proposed method, and boxplot of estimated ATE from 10,000 iterations. The bias and RMSE were calculated as $bias=\bar{\hat{\tau}}-\tau_{0}$, $RMSE=\sqrt{\frac{1}{10000}\sum_{k=1}^{10000}\left(\hat{\tau}_{k}-\tau_{0}\right)^{2}}$, where $\bar{\hat{\tau}}=\frac{1}{10000}\sum_{k=1}^{10000}\hat{\tau}_{k}$, $\hat{\tau}_{k}$ is the estimate of each estimator and iteration, and $\tau_{0}$ ($=0.152$) is the true value of the ATE. The CP refers to the proportion of cases where the confidence interval or credible interval includes $\tau_{0}$. The AvL was calculated as the average length of the Confidence (Credible) Interval over 10,000–time iterations. Bias and RMSE ratio (BR/RR) are calculated as the ratio compared to the bias/RMSE of our proposed method.
To evaluate the performance metrics of the BLPS and BRCAL, the posterior mean was used to assess the bias, RMSE, and BR/RR. For CP and AvL, the posterior distribution was used.
4.4 Simulation Results
The results are summarized in Tables 13 and Figure 1, with additional results provided in Appendix B. In the main manuscript, we only discuss the results under “Confounders” situations.
In small sample situations, all methods exhibit small bias in scenarios (a)-–(d). However, in scenarios (e) and (f), where the true propensity score model is complex, all methods—especially RCAL—yield large bias. This bias improves in moderate and large sample situations. The CBPS and BRCAL show similar performance, and notably, both methods exhibit the smallest bias in scenarios (e) and (f).
Across all scenarios (a)(f), the RMSE values are nearly identical for all methods. To some extent, the RCAL performs slightly better in small sample situations.
The 95% coverage probability (CP) of the BRCAL nearly achieves the nominal level and is higher than that of other methods in scenarios (a)(d). In scenario (e), the BRCAL slightly underestimates the 95% CI, but similar tendencies are observed in other methods. In small samples, all methods except the BRCAL and BLPS tend to underestimate the 95% CI in many scenarios. While the BLPS achieves the nominal level, it is overly conservative.
The BRCAL, our proposed Bayesian method, can be regarded as the Bayesian counterpart of the RCAL. However, its performance closely resembles that of the CBPS, particularly in large samples. This similarity may be explained by covariate balancing. For more details, see Appendix B.3.
To summarize, our proposed method demonstrates superior or at least comparable performance to other methods. Notably, it provides more accurate causal effect estimates and reliable inference, particularly in small samples. Covariate balancing diagnostics using standardized mean differences (SMDs) and sensitivity analysis results for the prior distribution of $\lambda$ are presented in Appendix B. Based on these analyses, we conclude that our proposed method achieves effective covariate balancing (similar to the CBPS) and remains robust to the choice of the prior distribution for $\lambda$.
[Tables 13 and Figure 1 here.]
5 Illustrative Real Data Example
In this section, we provide an illustrative example of a real data analysis conducted on the cohort study for male civil servants in London (Whitehall dataset)[29]. Specifically, we examined the causal effects of smoking on all-cause mortality. To estimate these effects, we considered three confounders: age, plasma cholesterol concentration, and Civil Service grade. To account for the binary exposure, we coded the original data with $A_{i}=0$ for non-smokers, and $A_{i}=1$ otherwise. The study included 17,260 subjects, of which 7,157 (41.5%) were smokers. There were also 1,670 deaths (9.7%). For more details, see Appendix C.
The summary results of the causal effects for each method are as follows: The crude mean difference is $6.86$ $(5.97,7.75)$ (point estimate (95% CI)), the CBPS result is $5.62$ $(5.58,5.68)$, and the RCAL result is $5.71$ $(4.81,6.62)$. The result of our proposed method is $5.65$ $(4.19,7.12)$ (posterior mean (95% CI)). Note that the specific sampling schemes are described in Section 3.2, and prior distributions of parameters are $\lambda\sim Gam(0.01,0.1)$, $\bm{\alpha}\sim DoubleExponential(0,1/\lambda)$, and $\theta_{k}\sim N(0,100^{2})$ ($k=0,1$). The point estimates are similar for the CBPS, RCAL, and BRCAL. The 95% CI for the BRCAL is the widest among the three methods. Given that there are 17,260 subjects, the CI is considered too wide. However, the CI is expected to provide conservative results, based on the findings from the simulation experiments.
6 Discussions and Conclusions
In this manuscript, we introduce a novel Bayesian-based IPW estimator that incorporates a covariate balancing condition. Using the general Bayes framework, which is gaining interest in Bayesian analysis, our proposed method does not require consideration of the full likelihood. Accordingly, it only necessitates minimal moment conditions, as is common in causal inference contexts. Our approach demonstrates reasonable performance, especially in small sample scenarios, as shown by various simulations and real data analyses.
To the best of our knowledge, this is the first method that applies the general Bayes framework to causal inference with a covariate balancing condition. As described in Section 3, the learning rate $\omega$ is determined by PCIC. This method proves effective, as evidenced by the 95% CI aligning well with the nominal probability. However, relying on PCIC is an ad-hoc strategy lacking theoretical justification. In future work, we aim to develop a consensus method that is theoretically grounded.
In our proposed framework, as mentioned in section 3.2, the posterior sampling scheme is conducted in two stages which allows us to estimate causal effects as in standard causal inference contexts [13]. The inference is justified from a Bayesian perspective [31], provided the second step is likelihood-based. Bissiri et al. [11] note that the general Bayesian procedure ensures valid Bayesian inference by meeting coherence and rationality [12]. Future work should explore further justification when applying the general Bayesian framework for the second step.
For future work, applying our approach to sequential treatment settings is crucial for estimating the ATE in situations with variable doses. While some methods incorporating covariate balancing conditions have been proposed [33, 34], a loss function akin to that described by Tan [10] has yet to be developed. Going forward, deriving such a loss function and considering a Bayesian estimation strategy following the same approach as presented in this manuscript will be essential. In sequential treatment settings, concerns about missing data, particularly in scenarios with extended follow-up periods, arise. In Appendix D, we consider the situations where one covariate is missing in point treatment scenarios. However, these methods cannot be directly applied to sequential settings. Xu et al. [35] introduce the Bayesian Additive Regression Tree method for imputing missing data as a Bayesian approach. This method, which is potentially adaptable to our proposed and future methodologies, could enable the construction of a comprehensive Bayesian method that effectively handles missing covariate data.
Author contributions: All authors have accepted responsibility for the entire content of this paper and approved its submission.
Acknowledgements: We would like to express our gratitude to the editors and reviewers for their useful comments, which have contributed greatly towards improving the manuscript. We would also like to thank Editage (www.editage.com) for the English language editing. All the authors accept responsibility for the content of this manuscript and approve its submission.
Data availability statement: Data sharing is not applicable to this article as no new data were created or analyzed in this study. The simulation analysis programs are available at the following URL:
  • https://github.com/t-momozaki/BRCAL
Appendix A Mathematical Calculations
A.1 Derivation of Loss Functions
First, we consider the loss function for the propensity score (2.2). Specifically, we consider the first derivatives respect to $\bm{\alpha}$:
$\frac{\partial}{\partial\bm{\alpha}}\ell_{A}(\bm{\alpha})$ $=\sum_{i=1}^{n}\left[-A_{i}\exp\left\{-\bm{\alpha}^{\top}g(\bm{X}_{i})\right\}g(\bm{X}_{i})+(1-A_{i})g(\bm{X}_{i})\right.$
$\hskip 14.22636pt\left.+(1-A_{i})\exp\left\{\bm{\alpha}^{\top}g(\bm{X}_{i})\right\}g(\bm{X}_{i})-A_{i}g(\bm{X}_{i})\right]$
$=\sum_{i=1}^{n}\left[-A_{i}\left\{1+\exp\left\{-\bm{\alpha}^{\top}g(\bm{X}_{i})\right\}\right\}+(1-A_{i})\left\{1+\exp\left\{\bm{\alpha}^{\top}g(\bm{X}_{i})\right\}\right\}\right]g(\bm{X}_{i})$
$=\sum_{i=1}^{n}\left[-\frac{A_{i}}{e(\bm{X}_{i};\bm{\alpha})}+\frac{1-A_{i}}{1-e(\bm{X}_{i};\bm{\alpha})}\right]g(\bm{X}_{i}),$
where we use the relationship from the propensity score model (i.e., logistic regression model) for the final equality. Additionally, for the penalty term (2.3),
$\frac{\partial}{\partial\bm{\alpha}}\lambda\|\bm{\alpha}\|_{1}=\lambda\times sgn(\bm{\alpha})=\left\{\begin{array}[]{lc}\lambda&\alpha_{\ell}>0\\ -\lambda&\alpha_{\ell}<0\\ \left[-\lambda,\lambda\right]&\alpha_{\ell}=0\\ \end{array}\right.$
From the results, the inequality (2.1) can be obtained.
Next, we consider the loss functions for the causal effect; here, we only consider the loss function (2.6) (i.e., for the treatment group). Considering the first derivatives respect to $\theta_{1}$,
$\frac{\partial}{\partial\theta_{1}}\ell_{Y_{1}}(\theta_{1},\bm{\alpha})=-2\sum_{i=1}^{n}\frac{A_{i}}{e_{i}(\bm{\alpha})}\left(Y_{i}-\theta_{1}\right).$
Therefore, a part of IPW estimator (2.4) can be obtained.
A.2 Derivation of Posterior Distributions
The posterior distributions for $\theta_{k}$ are derived as follows:
$p(\theta_{k}\mid\bm{Y},\bm{X},\bm{A},\bm{\alpha})$ $\propto\exp\left\{-\frac{\tau_{k}}{2}(\theta_{k}-\mu_{k})^{2}\right\}\exp\left\{-\omega\sum_{i=1}^{n}\frac{A_{i}^{k}(1-A_{i})^{1-k}}{e_{i}(\bm{\alpha})^{k}\{1-e_{i}(\bm{\alpha})\}^{1-k}}(Y_{i}-\theta_{k})^{2}\right\}$
$=\exp\left\{-\frac{\tau_{k}}{2}(\theta_{k}-\mu_{k})^{2}\right\}\exp\left\{-\frac{\omega}{2}\sum_{i=1}^{n}s_{ki}(Y_{i}-\theta_{k})^{2}\right\}$
$=\exp\left\{-\frac{1}{2}\left[\tau_{k}(\theta_{k}^{2}-2\mu_{k}\theta_{k}+\mu_{k}^{2})-\omega\sum_{i=1}^{n}s_{ki}(Y_{i}^{2}-2\theta_{k}Y_{i}+\theta_{k}^{2})\right]\right\}$
$\propto\exp\left\{-\frac{1}{2}\left[\left(\tau_{k}+\omega\sum_{i=1}^{n}s_{ki}\right)\theta_{k}^{2}-2\left(\tau_{k}\mu_{k}+\omega\sum_{i=1}^{n}s_{ki}Y_{i}\right)\theta_{k}\right]\right\}$
$\propto\exp\left\{-\frac{\tau_{k}+\omega\sum_{i=1}^{n}s_{ki}}{2}\left(\theta_{k}-\frac{\tau_{k}\mu_{k}+\omega\sum_{i=1}^{n}s_{ki}Y_{i}}{\tau_{k}+\omega\sum_{i=1}^{n}s_{ki}}\right)^{2}\right\}$
$=\exp\left\{-\frac{\tilde{\tau}_{k}}{2}(\theta_{k}-\tilde{\mu}_{k})^{2}\ \right\},$
where $\propto$” denotes proportionality with respect to $\theta_{k}$, and
Appendix B Additional Materials for Simulation Experiments
B.1 Details for Data-generating Mechanism
For each situation, the linear predictor $h_{A}\left(\bm{X}_{i}\right)$ of the true propensity score model was set as follows:
@@ B.1 Details for Data-generating Mechanism
B.2 Additional Results for Simulation Experiments
@@
The BRCAL and CBPS exhibit almost the same value of bias and RMSE in almost all scenarios. In contrast, the RCAL exhibits large, unstable bias compared to other methods. The results of all methods become consistent in large samples.
Table B.1: Summary of estimated causal effects under the scenario (a) and (b) (“Predictor” situation): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each scenario (in “Scenario” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning rate $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Scenario Method
@@
(a) Logit 0.183 6.426 0.930 0.232 0.835 1.023 0.134 3.928 0.945 0.150 0.956 1.010 0.105 2.764 0.946 0.107 0.789 1.005
CBPS 0.283 6.345 0.928 0.230 1.295 1.010 0.184 3.903 0.943 0.149 1.314 1.004 0.152 2.757 0.944 0.106 1.145 1.002
RCAL CV 0.556 5.751 0.932 0.208 2.541 0.915 0.337 3.779 0.919 0.132 2.412 0.972 0.134 2.704 0.917 0.093 1.007 0.983
$\lambda=0.01$ 0.548 6.448 0.903 0.208 0.300 3.795 0.919 0.132 0.190 2.642 0.923 0.093
$\lambda=0.05$ 0.514 5.440 0.944 0.209 0.866 3.392 0.946 0.131 1.011 2.380 0.948 0.093
$\lambda=0.10$ 0.362 5.236 0.950 0.207 0.399 3.303 0.953 0.130 0.321 2.328 0.951 0.092
BLPS 0.217 6.771 0.969 0.286 0.142 3.999 0.973 0.177 0.113 2.788 0.973 0.125
BRCAL PCIC 0.219 6.284 0.930 0.229 Ref. Ref. 0.140 3.889 0.936 0.145 Ref. Ref. 0.133 2.751 0.936 0.103 Ref. Ref.
$\omega=0.2$ 0.110 5.959 1.000 0.621 0.107 3.805 1.000 0.392 0.005 2.720 1.000 0.277
$\omega=0.5$ 0.107 6.172 0.999 0.393 0.051 3.863 0.999 0.248 0.087 2.740 0.999 0.175
$\omega=1.0$ 0.198 6.258 0.973 0.278 0.113 3.883 0.974 0.176 0.118 2.749 0.975 0.124
$\omega=1.5$ 0.219 6.287 0.928 0.227 0.140 3.889 0.935 0.143 0.133 2.751 0.934 0.101
(b) Logit 0.130 11.230 0.874 0.289 0.720 1.057 0.680 7.998 0.915 0.226 1.070 1.129 0.914 6.297 0.930 0.184 1.067 1.195
CBPS 0.245 10.856 0.958 0.520 1.351 1.022 0.678 7.147 0.915 0.248 1.067 1.009 0.868 5.293 0.903 0.175 1.013 1.005
RCAL CV 1.494 6.198 0.920 0.211 8.239 0.583 0.655 5.477 0.843 0.133 1.031 0.773 0.408 4.987 0.476 0.947
$\lambda=0.01$ 6.944 20.351 0.672 10.920 2.498 11.433 0.615 0.147 1.404 7.265 0.625 0.093
$\lambda=0.05$ 1.430 8.058 0.869 0.211 0.460 4.314 0.901 0.135 0.214 2.872 0.906 0.095
$\lambda=0.10$ 1.601 5.389 0.948 0.210 1.329 3.354 0.953 0.132 1.380 2.371 0.950 0.093
BLPS 0.319 12.837 0.877 0.342 0.754 8.612 0.833 0.202 0.978 6.603 0.792 0.138
BRCAL PCIC 0.181 10.624 0.832 0.351 Ref. Ref. 0.635 7.085 0.777 0.196 Ref. Ref. 0.857 5.268 0.753 0.134 Ref. Ref.
$\omega=0.2$ 0.057 10.313 0.993 0.598 0.546 6.952 0.993 0.385 0.789 5.205 0.990 0.274
$\omega=0.5$ 0.134 10.644 0.924 0.376 0.614 7.070 0.915 0.243 0.840 5.256 0.901 0.173
$\omega=1.0$ 0.196 10.751 0.794 0.265 0.652 7.110 0.782 0.172 0.859 5.275 0.761 0.122
$\omega=1.5$ 0.205 10.786 0.705 0.217 0.654 7.121 0.686 0.140 0.859 5.281 0.666 0.100
@@
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[4]; RCAL: Calibrating method proposed by Tan[10]; BLPS: Bayesian method proposed by Saarela et al.[Sa2015]; BRCAL: Proposed method;
(a): “Small complexity and good overlap,” (b): “Small complexity and poor overlap.”
Table B.2: Summary of estimated causal effects under the scenario (c) and (d) (“Predictor” situation): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each scenario (in “Scenario” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Scenario Method